Given the formula, solve for the indicated variable.
A = P(1 + nr) for r
Thank You!

Answer by stanbon(75887)

Given the formula, solve for the indicated variable.
A = P(1 + nr) for r
A/P = 1 + nr
nr = [A/P]-1
---
r = [(A/P)-1]/n
---
r = [(A-P)]/[nP]
==================
Cheers,
Stan H.